OpenStudy (anonymous):

let x=6, then evalute the expression 5x^2 + 2x -1

OpenStudy (anonymous):

5(6)^2 + 2(6) - 1 So, do your powers first so it becomes: =5(36) + 2(6) - 1 Then Multiply: =180 + 12 - 1 Then Add/Subtract: =191